# Episodic Contextual Bandits with Knapsacks under Conversion Models

### v1

@@ 1 Introduction
Lastly, our work is also related to contextual online RL, studied in [FosterKQR23, LevyM23, DengCZL24, QianHSC24, FosterHQR24]. In these models, there is only one context per episode. The context is generated before the episode begins, and the rewards and transition probability kernels in each time step during the epsiode are all dependent on that context. This is different from our model, which has $H$ contexts for the $H$ time steps during an episode, and each the context at time step $h$ only appears after time step $h-1$ but before time step $h$. In addition, by harnessing the problem structure with resource allocation, we achieve a better dependence on $H$ then those works, though their models and ours are not comparable in general.
Notation. We denote $\mathbb{R},\mathbb{Z}$ as the sets of real numbers and integers respectively, and denote $\mathbb{R}_{>0},\mathbb{Z}_{>0}$ as the sets of positive real numbers and positive integers respectively. For $n\in\mathbb{Z}_{>0}$, denote $[n]=\{1,\ldots,n\}$. For any $q\in[0,1]$, denote $\text{Bern}(q)$ as the Bernoulli distribution with mean $q$. For an event ${\cal E}$, denote $\mathbf{1}({\cal E})$ as its indicator random variable, where $\mathbf{1}({\cal E})=1$ if event ${\cal E}$ happens, and $\mathbf{1}({\cal E})=0$ otherwise. For $a,b\in\mathbb{R}$, denote $a\wedge b=\min\{a,b\}$. Denote the size of a set ${\cal A}$ as $|{\cal A}|$.
@@ 2.1 Model
Dynamics. At the start of episode $t$, the DM is endowed with $B_{t}\in[LH]$ units of the resource. The DM only observes the value of $B_{t}$ after the end of episode $t-1$, but before the start of episode $t$. Denote $b_{h,t}$ as the amounts of the remaining resource at the start of the time step $(h,t)$ (observe that $b_{1,t}=B_{t}$). At the time step $(h,t)$, firstly a request with context $\theta_{h,t}\in\Theta$ appears. The context $\theta_{h,t}$ is distributed according to the latent probability distribution $\Lambda_{h}$. The contexts $(\theta_{h,t})^{H}_{h=1}\sim\prod^{H}_{h=1}\Lambda_{h}$ are independent but not necessarily identically distributed, since $\Lambda_{1},\ldots,\Lambda_{H}$ vary in general. After that, the DM selects an allocation decision $A_{h,t}\in{\cal A}(b_{h,t},\theta_{h,t})$, based on $\theta_{h,t}$, $b_{h,t}$ and the observations obtained prior to time step $(h,t)$. The restricted set ${\cal A}(b_{h,t},\theta_{h,t})$ ensures that the DM consumes at most $B_{t}$ resource units in episode $t$, and does not violate the resource constraint.
@@
$\rho(\theta,a)$ Pr(customer $\Pr(\text{customer buys})$ Pr(worker $\Pr(\text{worker accepts})$ Pr(DM wins the $\Pr(\text{DM wins the auction})$
$r(\theta,a)$ $a$ $1$ $\theta-a$
$d(\theta,a)$ $1$ $a$ $a$
@@ 2.2 Characterizing $\textsf{opt}_{1}\ldots,\textsf{opt}_{T}$ with Bellman Equations
and we set $V_{H+1}(b,\theta)=0$ for all $b,\theta$. Observe $\textsf{opt}_{t}=\sum_{\bar{\theta}\in\Theta}\Pr_{\theta_{1}\sim\Lambda_{1}}(\theta_{1}=\bar{\theta})\cdot V_{1}(B_{t},\bar{\theta})$. The set of Bellman equations (1) highlights the challenge with the large state space ($(\{0\}\cup[LH])\times\Theta$. The latent transition probability is of the form $\rho(\theta,a)\cdot\Pr_{\theta_{h+1}\sim\Lambda_{h+1}}(\theta_{h+1}=\bar{\theta})$, thus the number of latent parameters is linear in $H\cdot|{\cal A}|\cdot|\Theta|$, which is intractably large. A reason of the intractability is the term $\Pr_{\theta_{h+1}\sim\Lambda_{h+1}}(\theta_{h+1}=\bar{\theta})$, which is in general unstructured, and typical functional approximation techniques in RL doe not apply. Lastly, the above discussion assumes a finite or countably infinite $\Theta$, but we can handle an uncountably infinite (but measurable) $\Theta$ by replacing $\sum_{\bar{\theta}\in\Theta}\Pr_{\theta_{h+1}\sim\Lambda_{h+1}}(\theta_{h+1}=\bar{\theta})\cdot V_{h+1}(b^{\prime},\bar{\theta})$ with $\mathbb{E}_{\theta_{h+1}\sim\Lambda_{h+1}}[V_{h+1}(b^{\prime},\theta_{h+1})]$, which we denote as $U_{h+1}(b^{\prime})$.
@@ 3.1 Algorithm Design
We propose the online algorithm Mimic-Opt-DP, shown in Algorithm 1. We first explain the main steps, then discuss how Mimic-Opt-DP estimates the latent parameters. Mimic-Opt-DP involves two sets of estimates: (i) $\{\hat{U}_{h+1,t}\}_{h\in[H+1]}$: $\hat{U}_{h+1,t}(b)$ serves to estimate $U_{h+1}(b)=\mathbb{E}_{\theta_{h+1}\sim\Lambda_{h+1}}[V_{h+1}(b,\theta_{h+1})]$, the optimal expected reward from time step $h+1$ to $H$, when the DM has $b$ units of remaining resources at the start of time step $h+1$. (ii) $\text{UCB}_{t},\text{LCB}_{t}$, which serves to estimate $\rho$. The estimates $\hat{U}_{h+1,t}$, $\text{UCB}_{t},\text{LCB}_{t}$, constructed using data in episodes $11,\ldots,t-$ 1, are for choosing the action $A_{h,t}$ at time step $(h,t)$ (Line 7 in Algorithm 1). The optimization problem (3) serves to estimate the Bellman equation (1). To formulate (3), we replace the latent $\rho$ with its optimistic estimates, where we use upper confidence bounds (UCBs) for the reward earned and lower confidence bounds (LCBs) for the resource consumed. The replacement thus follows the principle of Optimism in Face of Uncertainty.
We construct $\{\hat{U}_{h,t}\}_{h\in[H+1]}$ in Optimistic-DP displayed in Algorithm 2. The construction, shown in Lines 3 to 5 in Algorithm 2, involves a backward induction that resembles the Bellman equations (1). We replace the latent parameters in (1) with their estimates similarly to (3). In this way, the action selection in (3) in fact mimics the way $\{\hat{U}_{h,t}\}_{h\in[H+1]}$ are constructed. The backward induction in Algorithm 2 requires estimating both the latent quantities $\rho$ and $\{\Lambda_{h}\}_{h\in[H]}$. At the end of episode $t-1$, we estimate $\rho$ and $\{\Lambda_{h}\}_{h\in[H]}$ using two disjoint datasets ${\cal D}_{t}$ and ${\cal S}_{t}$ respectively. Their disjointness facilitates our analysis. We use the data in episodes in ${\cal J}_{t}\subset[t-1]$ to estimate $\{\Lambda_{h}\}_{h\in[H]}$, and the data in episodes in $[t-1]\setminus{\cal J}_{t}$ to estimate $\rho$. Our default choice of $\{{\cal J}_{t}\}^{\infty}_{t=1}$ depends on an absolute constant $\alpha\in(0,1)$, which we take $\alpha=1/2$ by default. Let $n_{\alpha}\in\mathbb{Z}_{>0}$ satisfy $(1/\alpha)-1\leq n_{\alpha}<(2/\alpha)-1$. We define the default choice
@@ 3.2 Main Results
where $\tilde{O}(\cdot)$ hides a multiplicative factor of absolute constant times $\sqrt{\log(LHT/\delta)}$, and $O(\cdot)$ hides a multiplicative factor of absolute constant. In particular, the regret bound is independent of $|\Theta|$.
@@
The second term is due to estimating $\rho$ via the CB oracle ${\cal O}$. Recall that $[T]\setminus{\cal J}_{T+1}$ consists of the episodes when we invoke ${\cal O}$. Thus, $(\text{Loss})=\sum_{t\in[T]\setminus{\cal J}_{T+1}}\sum^{H}_{h=1}[\text{UCB}_{t}(\theta_{h,t},A_{h,t})-\text{LCB}_{t}(\theta_{h,t},A_{h,t})]$ reflects how fast the confidence radius under ${\cal O}$ shrinks, when we update the UCBs, LCBs episode by episode, and each episode generates $\leq H$ samples (Line 15 in Algorithm 1). As surveyed in Appendix A.2, existing research works on batched contextual bandits [RenZK22] imply that $(\text{Loss})=o(T)$ and independent of $|\Theta|$ for models such as contextual generalized linear bandit model [FilippiCGS10], contextual logistic bandit model [FauryACF2020] and contextual Gaussian process bandit model [KrauseO11].
To further illustrate, we specialize Theorem 3.1 on some of the applications in Table 1. The specialized regret bounds are displayed in the “$M=0,{\cal S}_{1}=\emptyset$” column in Table 2, where we consider the applications with dynamic pricing (see Appendix A.1.1) and first price auctions (see Appendix A.1.3). More applications can be found and their regret bounds are in Appendix A.1. In the first row on dynamic pricing with generalized linear bandit model, the parameters $\ell_{f},\kappa_{f}>0$ are bounding parameters related to a demand-curve function $f$ that models $\rho$, see Appendix A.1.1.
@@
gen. linear model $+r_{\text{max}}d\cdot\frac{\ell_{f}}{\kappa_{f}}\sqrt{HT}+r_{\text{max}}dH)$ $\tilde{O}(r_{\text{max}}d\cdot\frac{\ell_{f}}{\kappa_{f}}\sqrt{HT}+r_{\text{max}}dH)$
@@
identical items $+r_{\text{max}}L\sqrt{|{\cal A}|HT}+r_{\text{max}}L|{\cal A}|H)$ $\tilde{O}(r_{\text{max}}L\sqrt{|{\cal A}|HT}+r_{\text{max}}L|{\cal A}|H)$
@@
The event ${\cal F}_{t,\delta^{\prime}}$ is defined in Lemma 3.4, the event ${\cal G}_{t,\delta_{t}}$ is defined in (6). The event ~$\tilde{\cal E}_{h,t,\delta^{\prime}}$, which satisfies $\Pr(\tilde{\cal E}_{h,t,\delta^{\prime}})\geq 1-\delta^{\prime}/H$, is defined in (54) in Appendix B.3. The term $(\star_{t})$ is a random variable with the properties that $\mathbb{E}[(\star_{t})]=0$, and $\Pr(\sum^{T}_{t=1}(\star_{t})\leq 2r_{\text{max}}(2L+1)\sqrt{HT\log(2/\delta^{\prime\prime})})\geq 1-3\delta^{\prime\prime}$ for any $\delta^{\prime\prime}\in(0,1)$.
@@
The upper bound (9) is because consecutive elements in (2) in ${\cal J}_{T+1}$ are at most $(2/\alpha)-1$ apart, and $|{\cal J}_{T+1}|\leq\alpha T$. To show $\text{Reg}(T)=o(T)$, it suffices to show $\sum^{T}_{t=1}\sum^{H}_{h=1}[\text{UCB}_{t}(\theta_{h,t},A_{h,t})-\text{LCB}_{t}(\theta_{h,t},A_{h,t})]=o(T)$. While such a bound appears similar to existing results on UCB-based algorithms, we caution that in our algorithm we only use the data in episodes indexed by $[t-1]\setminus{\cal J}_{t}$ to construct $\text{UCB}_{t},\text{LCB}_{t}$, different from existing results where all data in episodes/rounds indexed by $[t-1]$ are used to construct the upper or lower confidence bounds at $t$. We reconcile the connection to existing results on UCB by the following:
@@ 3.4 Benefits of Unlabeled Feature Data
Mimic-Opt-DP benefits from the availability of ${\cal S}=\{(\theta_{h,-m})_{h\in[H]}\}_{m\in[M]}$, and enjoys improved regret bounds in some settings. Consider the following index sets $\{\tilde{\cal J}_{t}\}^{\infty}_{t=1}$ parameterized by an absolute constant $\alpha\in(0,1)$ and a positive integer $n_{\alpha}$ satisfying $11/\alpha\leq n_{\alpha}<(2/\alpha)-$ 1:
@@
Corollary 3.7 has an interesting implication on efficiently approximating the optimum opt in an episode, which has a fixed but arbitrary starting inventory $\bar{B}\in[LH]$. Let’s assume (only in this paragraph) that the DM knows $\rho$, so that he can set $\text{UCB}_{t}=\text{LCB}_{t}=\rho$, and the DM possesses the unlabeled feature data ${\cal S}=\{(\theta_{h,-m})_{h\in[H]}\}_{m\in[M]}$, but still does not know $\{\Lambda_{h}\}^{H}_{h=1}$. Approximating opt is still challenging due to the intractability of the Bellman equations (1). Consider an instance ${\cal I}$ with $M$ episode, where $B_{t}=\bar{B}$ for all $t\in[M]$. Let’s run Mimic-Opt-DP on instance ${\cal I}$ with input ${\cal S}_{1}={\cal S}$ and CB oracle knowing $\rho$. Corollary 3.7 guarantees that $\textsf{opt}-\frac{1}{M}\sum^{M}_{t=1}\sum_{h=1}^{H}R_{h,t}\leq\tilde{O}(r_{\text{max}}LH/\sqrt{M})$ with probability $\geq 1-9\delta$. Next, the Azuma-Hoeffding inequality ensures the latter inequality in $\textsf{opt}-\frac{1}{M}\sum^{M}_{t=1}\sum_{h=1}^{H}R_{h,t}\geq\mathbb{E}[\frac{1}{M}\sum^{M}_{t=1}\sum_{h=1}^{H}R_{h,t}]-\frac{1}{M}\sum^{M}_{t=1}\sum_{h=1}^{H}R_{h,t}\geq-\tilde{O}(r_{\text{max}}H/\sqrt{M})$ holds with probability $1-\delta$. Altogether, we have the approximation guarantee $|\textsf{opt}-\frac{1}{M}\sum^{M}_{t=1}\sum_{h=1}^{H}R_{h,t}|\leq\tilde{O}(r_{\text{max}}LH/\sqrt{M})$ which holds with probability $\geq 1-10\delta$, and the $\tilde{O}(\cdot)$ hides a multiplicative factor of absolute constant times $\sqrt{\log(LHM/\delta)}$ independent of $|\Theta|$.

### v2

@@ 1 Introduction
Lastly, our work is also related to contextual online RL, studied in [FosterKQR23, LevyM23, DengCZL24, QianHSC24, FosterHQR24]. In these models, there is only one context per episode. The context is generated before the episode begins, and the rewards and transition probability kernels in each time step during the epsiode are all dependent on that context. This is different from our model, which has $H$ contexts for the $H$ time steps during an episode, and each the context at time step $h$ only appears after time step h1h-1 but before time step $h$. In addition, by harnessing the problem structure with resource allocation, we achieve a better dependence on $H$ then those works, though their models and ours are not comparable in general.
[…]
@@ 2.1 Model
Dynamics. At the start of episode $t$, the DM is endowed with $B_{t}\in[LH]$ units of the resource. The DM only observes the value of $B_{t}$ after the end of episode t1t-1, but before the start of episode $t$. Denote $b_{h,t}$ as the amounts of the remaining resource at the start of the time step $(h,t)$ (observe that $b_{1,t}=B_{t}$). At the time step $(h,t)$, firstly a request with context $\theta_{h,t}\in\Theta$ appears. The context $\theta_{h,t}$ is distributed according to the latent probability distribution $\Lambda_{h}$. The contexts $(\theta_{h,t})^{H}_{h=1}\sim\prod^{H}_{h=1}\Lambda_{h}$ are independent but not necessarily identically distributed, since $\Lambda_{1},\ldots,\Lambda_{H}$ vary in general. After that, the DM selects an allocation decision $A_{h,t}\in{\cal A}(b_{h,t},\theta_{h,t})$, based on $\theta_{h,t}$, $b_{h,t}$ and the observations obtained prior to time step $(h,t)$. The restricted set ${\cal A}(b_{h,t},\theta_{h,t})$ ensures that the DM consumes at most $B_{t}$ resource units in episode $t$, and does not violate the resource constraint.
@@
$\rho(\theta,a)$ $\Pr(\text{customer buys})$ $\Pr(\text{worker accepts})$ $\Pr(\text{DM wins the auction})$
$r(\theta,a)$ $a$ 11 $\theta-a$
$d(\theta,a)$ 11 $a$ $a$
@@ 2.2 Characterizing $\textsf{opt}_{1}\ldots,\textsf{opt}_{T}$ with Bellman Equations
and we set $V_{H+1}(b,\theta)=0$ for all $b,\theta$. Observe $\textsf{opt}_{t}=\sum_{\bar{\theta}\in\Theta}\Pr_{\theta_{1}\sim\Lambda_{1}}(\theta_{1}=\bar{\theta})\cdot V_{1}(B_{t},\bar{\theta})$. The set of Bellman equations (1) highlights the challenge with the large state space $(\{0\}\cup[LH])\times\Theta$. The latent transition probability is of the form $\rho(\theta,a)\cdot\Pr_{\theta_{h+1}\sim\Lambda_{h+1}}(\theta_{h+1}=\bar{\theta})$, thus the number of latent parameters is linear in $H\cdot|{\cal A}|\cdot|\Theta|$, which is intractably large. A reason of the intractability is the term $\Pr_{\theta_{h+1}\sim\Lambda_{h+1}}(\theta_{h+1}=\bar{\theta})$, which is in general unstructured, and typical functional approximation techniques in RL doe not apply. Lastly, the above discussion assumes a finite or countably infinite $\Theta$, but we can handle an uncountably infinite (but measurable) $\Theta$ by replacing $\sum_{\bar{\theta}\in\Theta}\Pr_{\theta_{h+1}\sim\Lambda_{h+1}}(\theta_{h+1}=\bar{\theta})\cdot V_{h+1}(b^{\prime},\bar{\theta})$ with $\mathbb{E}_{\theta_{h+1}\sim\Lambda_{h+1}}[V_{h+1}(b^{\prime},\theta_{h+1})]$, which we denote as $U_{h+1}(b^{\prime})$.
@@ 3.1 Algorithm Design
We propose the online algorithm Mimic-Opt-DP, shown in Algorithm 1. We first explain the main steps, then discuss how Mimic-Opt-DP estimates the latent parameters. Mimic-Opt-DP involves two sets of estimates: (i) $\{\hat{U}_{h+1,t}\}_{h\in[H+1]}$: $\hat{U}_{h+1,t}(b)$ serves to estimate $U_{h+1}(b)=\mathbb{E}_{\theta_{h+1}\sim\Lambda_{h+1}}[V_{h+1}(b,\theta_{h+1})]$, the optimal expected reward from time step $h+1$ to $H$, when the DM has $b$ units of remaining resources at the start of time step $h+1$. (ii) $\text{UCB}_{t},\text{LCB}_{t}$, which serves to estimate $\rho$. The estimates $\hat{U}_{h+1,t}$, $\text{UCB}_{t},\text{LCB}_{t}$, constructed using data in episodes $1,\ldots,t-1$, are for choosing the action $A_{h,t}$ at time step $(h,t)$ (Line 7 in Algorithm 1). The optimization problem (3) serves to estimate the Bellman equation (1). To formulate (3), we replace the latent $\rho$ with its optimistic estimates, where we use upper confidence bounds (UCBs) for the reward earned and lower confidence bounds (LCBs) for the resource consumed. The replacement thus follows the principle of Optimism in Face of Uncertainty.
We construct $\{\hat{U}_{h,t}\}_{h\in[H+1]}$ in Optimistic-DP displayed in Algorithm 2. The construction, shown in Lines 3 to 5 in Algorithm 2, involves a backward induction that resembles the Bellman equations (1). We replace the latent parameters in (1) with their estimates similarly to (3). In this way, the action selection in (3) in fact mimics the way $\{\hat{U}_{h,t}\}_{h\in[H+1]}$ are constructed. The backward induction in Algorithm 2 requires estimating both the latent quantities $\rho$ and $\{\Lambda_{h}\}_{h\in[H]}$. At the end of episode t1t-1, we estimate $\rho$ and $\{\Lambda_{h}\}_{h\in[H]}$ using two disjoint datasets ${\cal D}_{t}$ and ${\cal S}_{t}$ respectively. Their disjointness facilitates our analysis. We use the data in episodes in ${\cal J}_{t}\subset[t-1]$ to estimate $\{\Lambda_{h}\}_{h\in[H]}$, and the data in episodes in $t-1]\setminus{\cal J}_{t}$ to estimate $\rho$. Our default choice of $\{{\cal J}_{t}\}^{\infty}_{t=1}$ depends on an absolute constant $\alpha\in(0,1)$, which we take $\alpha=1/2$ by default. Let $n_{\alpha}\in\mathbb{Z}_{>0}$ satisfy $(1/\alpha)-1\leq n_{\alpha}<(2/\alpha)-1$. We define the default choice
@@ 3.2 Main Results
[…]
@@
The second term is due to estimating $\rho$ via the CB oracle $\cal O$}. Recall that $T]\setminus{\cal J}_{T+1}$ consists of the episodes when we invoke $\cal O$}. Thus, $(\text{Loss})=\sum_{t\in[T]\setminus{\cal J}_{T+1}}\sum^{H}_{h=1}[\text{UCB}_{t}(\theta_{h,t},A_{h,t})-\text{LCB}_{t}(\theta_{h,t},A_{h,t})]$ reflects how fast the confidence radius under $\cal O$} shrinks, when we update the UCBs, LCBs episode by episode, and each episode generates $\leq H$ samples (Line 15 in Algorithm 1). As surveyed in Appendix A.2, existing research works on batched contextual bandits [RenZK22] imply that $(\text{Loss})=o(T)$ and independent of $|\Theta|$ for models such as contextual generalized linear bandit model [FilippiCGS10], contextual logistic bandit model [FauryACF2020] and contextual Gaussian process bandit model [KrauseO11].
To further illustrate, we specialize Theorem 3.1 on some of the applications in Table 1. The specialized regret bounds are displayed in the $M=0,{\cal S}_{1}=\emptyset$” column in Table 2, where we consider the applications with dynamic pricing (see Appendix A.1.1) and first price auctions (see Appendix A.1.3). More applications can be found and their regret bounds are in Appendix A.1. In the first row on dynamic pricing with generalized linear bandit model, the parameters $\ell_{f},\kappa_{f}>0$ are bounding parameters related to a demand-curve function $f$ that models $\rho$, see Appendix A.1.1.
@@
gen. linear model $r_{\text{max}}d\cdot\frac{\ell_{f}}{\kappa_{f}}\sqrt{HT}+r_{\text{max}}dH)$ $\tilde{O}(r_{\text{max}}d\cdot\frac{\ell_{f}}{\kappa_{f}}\sqrt{HT}+r_{\text{max}}dH)$
@@
identical items $r_{\text{max}}L\sqrt{|{\cal A}|HT}+r_{\text{max}}L|{\cal A}|H)$ $\tilde{O}(r_{\text{max}}L\sqrt{|{\cal A}|HT}+r_{\text{max}}L|{\cal A}|H)$
@@
The event ${\cal F}_{t,\delta^{\prime}}$ is defined in Lemma 3.4, the event ${\cal G}_{t,\delta_{t}}$ is defined in (6). The event $\tilde{\cal E}_{h,t,\delta^{\prime}}$, which satisfies $\Pr(\tilde{\cal E}_{h,t,\delta^{\prime}})\geq 1-\delta^{\prime}/H$, is defined in (54) in Appendix B.3. The term $(\star_{t})$ is a random variable with the properties that $\mathbb{E}[(\star_{t})]=0$, and $\Pr(\sum^{T}_{t=1}(\star_{t})\leq 2r_{\text{max}}(2L+1)\sqrt{HT\log(2/\delta^{\prime\prime})})\geq 1-3\delta^{\prime\prime}$ for any $\delta^{\prime\prime}\in(0,1)$.
@@
The upper bound (9) is because consecutive elements in (2) in ${\cal J}_{T+1}$ are at most $(2/\alpha)-1$ apart, and $|{\cal J}_{T+1}|\leq\alpha T$. To show $\text{Reg}(T)=o(T)$, it suffices to show $\sum^{T}_{t=1}\sum^{H}_{h=1}[\text{UCB}_{t}(\theta_{h,t},A_{h,t})-\text{LCB}_{t}(\theta_{h,t},A_{h,t})]=o(T)$. While such a bound appears similar to existing results on UCB-based algorithms, we caution that in our algorithm we only use the data in episodes indexed by $t-1]\setminus{\cal J}_{t}$ to construct $\text{UCB}_{t},\text{LCB}_{t}$, different from existing results where all data in episodes/rounds indexed by [t1][t-1] are used to construct the upper or lower confidence bounds at $t$. We reconcile the connection to existing results on UCB by the following:
@@ 3.4 Benefits of Unlabeled Feature Data
Mimic-Opt-DP benefits from the availability of ${\cal S}=\{(\theta_{h,-m})_{h\in[H]}\}_{m\in[M]}$, and enjoys improved regret bounds in some settings. Consider the following index sets $\{\tilde{\cal J}_{t}\}^{\infty}_{t=1}$ parameterized by an absolute constant $\alpha\in(0,1)$ and a positive integer $n_{\alpha}$ satisfying $1/\alpha\leq n_{\alpha}<(2/\alpha)-1$: